# Christian Ghiglino    Alastair Langtry

### v4

@@ Abstract
This paper adapts ideas from social identity theory to set out a new framework for modelling conspicuous consumption. Agents derive status from their own conspicuous consumption and from belonging to an identity group with high conspicuous consumption. Importantly, these two sources of status are substitutes. Agents also feel pressure to conform with their neighbours in a network. This framework can rationalise a set of seemingly conflicting stylised facts about conspicuous consumption that are currently explained by different families of models. In addition, our model delivers new testable predictions regarding the effect of network structure and income inequality on conspicuous consumption.
@@
Keywords:      Social identity, Keeping Up with the Joneses, networks, centrality,
conspicuous consumption, income inequality
Economists widely recognise that people’s consumption decisions respond to the consumption decisions of others. Existing literature documents two main stylised facts about these responses. First, an influential paper by Charles et al. (2009) documents that, holding their own permanent income fixed, people in the US consume less when average income of those of the same race and who live in the same state is higher. Khamis et al. (2012) find similar behaviour for caste/religious groups in India. Second, work by Kuhn et al. (2011) shows that people consume more when average income of their geographic neighbours increases. This is often called a “Keeping up with The Joneses” effect. On the face of it, these two stylised facts may appear to conflict. Plus, the classes of models typically used to rationalise each of them – signalling models for the Charles et al. (2009) result and social comparisons models for the Kuhn et al. (2011) result – cannot capture both simultaneously.
We present a simple model of social identity that explains both of these stylised facts as a single phenomenon. In it, people derive status directly from consuming visible goods themselves, and from belonging to an identity that consumes lots of visible goods. They may also get intrinsic benefits from their own consumption, but adding this has no impact on the insights or the outcomes. In our model, identity group status depends on average consumption of visible goods by members of the identity group, and we assume there is some substitutability between status from each of these two sources. People also feel pressure to conform to some benchmark for their identity, and we assume it depends only on a person’s neighbours rather than their entire identity group. These two features are both core components of social identity theory (Costa-Font and Cowell, 2015). Our model is also able to speak to the first stylised fact without relying on the characteristics in the lowest part of the distribution – which is needed in signalling models – aligning our model and its predictions more closely with available data.
We also explain two further stylised features of consumption decisions. Third, Kaus (2013) finds that richer people are more sensitive to changes in the income of others. Our model predicts exactly this behaviour. [^2] A large part of the contribution in Kaus (2013) is to extend the analysis in Charles et al. (2009) to a lower income country – South Africa. He replicates the findings for Black South Africans (who make up the large majority of the South African population), but not for White South Africans. But the finding regarding sensitivity is novel to Kaus (2013). Fourth, Charles et al. (2009) find that higher dispersion of income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks. Our model is flexible enough to capture these contrasting behaviours. It can do so because it finds that the overall impact of a change in inequality depends on the network structure. Therefore we can explain both of these findings if Blacks and Whites have different network structures. We believe explaining the fourth stylised fact in this way is a useful contribution because it is able to capture contrasting behaviours without relying on differences in preferences across racial groups. This is something that existing signalling models can struggle with.
@@
On the way to explaining these four stylised facts, we provide a simple characterisation of behaviour. Equilibrium consumption depends critically on a generalised version of Bonacich centrality (Katz, 1953; Bonacich, 1987) – a well-known measure of network centrality. But here, both an agent’s own centrality and the average centrality among her identity group are important. In line with existing work, centrality is increasing in the strength of network connections, in an agent’s own income and in the income of those who share her identity. Related to this, we find that stronger homophily – when people are more tightly connected to those like themselves – increases consumption. This is because being more connected to people who share her identity increases the pressure an agent feels to conform – which pushes up consumption.
We also study the effects of changing inequality. We consider a setting with a number of separate communities, and examine the inequality between different communities and within a given community separately. The impact of redistributing across different communities depends critically on the relative densities of network connections in the affected communities. This highlights the role that network connections play in mediating the impact of inequality. When considering the impact of redistribution within a community, we impose a regularity condition on the network to simplify the analysis. Subject to this condition, the impact depends on whether the redistribution is progressive or regressive. A progressive redistribution increases overall consumption, while a regressive redistribution does the opposite. This lines up with the standard intuition that poorer agents’ consumption choices are more responsive to changes in incomes.
The key novelty compared to existing models of conspicuous consumption is that we have two sources of status – own consumption and average consumption of the identity group – and that they are imperfect substitutes. That status comes from one’s identity group is a core idea in the social identity theory literature. But as in existing models of social comparisons, agents still feel pressure to keep up with their friends – creating a strategic complementarity between a person’s own consumption and that of her neighbours. [^3] One difference to existing models of social comparisons is that agents in our model feel pressure to conform to their friends – they dislike falling behind, but also dislike getting too far ahead – rather than feeling pressure to get ahead. But this is not important for the results. So when one agent’s consumption changes, whether this leads some other agent to raise or lower her consumption in response depends on how closely connected the two are in the network.
To see how this explains the two main stylised facts, suppose that incomes increase for some people. These people increase their consumption of visible goods – the traditional direct response to an increase in own income. Now consider a single person whose income did not change. The change in other people’s income creates two spillover effects. First, it increases the status benefits for the single person, by increasing the average consumption of those who share her identity. This then pushes her to consume fewer visible goods – because the status from her identity and the status from her consumption of visible goods are substitutes. Second, it increases her neighbours’ consumption, which increases the pressure to consume visible goods through the dissonance costs. This pushes her to consume more visible goods.
@@
Which of the two forces wins out depends on how closely connected the single person is to the people whose incomes increase. This is precisely what delivers both of the stylised facts. When the people whose incomes increase are close neighbours (as in Kuhn et al. (2011) and Khamis et al. (2012)) we should expect the second force to win out – so she consumes more visible goods. In contrast, when these people are not closely connected to her (as in Charles et al. (2009)) we should expect the first force to win out – delivering the seemingly counter-intuitive result that she consumes fewer visible goods.
@@
The rest of the paper is organised in the usual way. Section 1 briefly reviews related literature. Section 2 presents the model. Section 3 presents the results. Section 4 shows how our model explains the stylised facts. Then Section 5 examines the impact of inequality. Section 6 concludes.
@@ 1 Literature
Our paper sits within a large literature on interpersonal comparisons and conspicuous consumption. [^4] It spans disciplines, including marketing, sociology, social psychology (see Dubois and Ordabayeva (2017) and Dubois et al. (2021) for recent reviews), and economics (where it dates at least to Veblen (1899), with other early contributions from Duesenberry (1949); Festinger (1954); Frank (1985b, a) and Van de Stadt et al. (1985)). Bursztyn and Jensen (2017) provide an interesting survey of the wider topic of social pressure. Within economics, there are two broad views on why people engage in conspicuous consumption. The first assumes that people care about their income or wealth relative to others – and conspicuous consumption is used as a signal of high income/wealth. Charles et al. (2009) and Khamis et al. (2012) use this signalling approach to explain the first stylised fact – when others get richer, a given agent moves down the income ranking, and so has fewer people she needs to signal she is richer than (their model follows Glazer and Konrad (1996) very closely). However, with no explicit notion of ‘neighbours’, these models struggle to explain the second stylised fact – that people consume more when their neighbours’ incomes rise. Additionally, because agents are consuming to differentiate themselves from people poorer than themselves, everyone’s consumption decisions can be sensitive to the income of those at the very bottom of the distribution.
@@
The second view assumes that people make social comparisons direction over consumption – so people gain status from conspicuous consumption itself. The core idea is that people like high consumption relative to others, not just in absolute terms – often called a “Keeping up with the Joneses” motive. It is perhaps the more dominant of the two views. Earlier work assumed people either care about their consumption relative to the population average, or to their rank within the population. Important contributions to these two approaches include Abel (1990); Clark and Oswald (1996); Ljungqvist and Uhlig (2000), and (Frank, 1985a, b, 2001; Hopkins and Kornienko, 2004) respectively. [^5] More recent contributions include Luttmer (2005); Heffetz (2011); Frank et al. (2014); Drechsel-Grau and Schmid (2014); Alvarez-Cuadrado et al. (2016); Jinkins (2016); Bertrand and Morse (2016),De Giorgi et al. (2020) and Hopkins (2023) Our paper is most closely related to a recent strand within this literature, started by Ghiglino and Goyal (2010), that uses explicit networks to allow for very rich patterns of social comparisons (more recent contributions to this strand include Immorlica et al. (2017); Langtry (2022) and Bramoullé and Ghiglino (2022)). These ‘Keeping up with the Joneses’ models successfully explain the second stylised fact. But they cannot explain the first. This is because they embed a desire to ‘Keep up with the Joneses’, which makes agents’ consumption choices strategic complements – an increase in some people’s income must (weakly) increase everyone’s consumption.
@@
What our paper adds is the ability to explain a set of so-far difficult to reconcile stylised facts in a single framework. It does so by building a new type of model – one motivated by social identity, rather than signalling or ‘Keeping up with the Joneses’ motives. In doing so, we also relate to a strand of literature that examines the link between inequality and conspicuous consumption. By and large, formal models provide ambiguous results on the nature of the link. In Charles et al. (2009), a regressive redistribution of income (from a poorer agent to a richer one) increases total conspicuous consumption only if consumption is convex in income – but this is not a consistent property of observed consumption data (Heffetz, 2011). [^6] Other versions of the signalling model make similar predictions (Ireland, 2001; Heffetz, 2011; Moav and and Neeman, 2012; Jinkins, 2016). Further, Bilancini and Boncinelli (2012) show that the relationship between conspicuous consumption and inequality can be sensitive to the choice of whether to use ordinal or cardinal utility (also see Hopkins (2023) for similar discussion).
@@
In line with this literature, we do not present an ironclad relationship between inequality and conspicuous consumption. But we do provide novel predictions regarding how the nature of the relationship depends on properties of the network. Notably, it allows us to explain our fourth stylised fact – that higher dispersion in income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks. It does so without an appeal to differences in underlying preferences across groups, but instead relies on well-documented differences in average income.
@@
Finally, our paper relates to the burgeoning literature on social identity theory within economics (see Costa-Font and Cowell (2015) for a survey). The core idea in these models is that agents are part of an ‘identity’ group, and gain some status from belonging to the group, but also bear some costs from deviating from some benchmark for that group (either real differences, or in the agents’ own perception of matters). This setup was formalised in seminal contributions from Akerlof and Kranton (2000) and Shayo (2009). [^7] A number of recent papers extend or apply Shayo’s approach; see for example, Klor and Shayo (2010); Gennaioli and Tabellini (2019); Grossman and Helpman (2021); Lindqvist and Östling (2013), and Shayo (2020). These papers allow people to choose their identity group, but keep each identity’s prototype (the benchmark for ‘normal’) exogenously fixed.
@@
Our main contribution here is to show how models of social identity can help resolve a puzzle regarding conspicuous consumption behaviour – something these models have not been used for to date. From a more technical standpoint, we also make the benchmark that agents find it costly to deviate from (often called a ‘prototype’ in the literature) endogenous. Aside from this, our utility formulation closely follows Shayo (2009). However, we abstract away from the choice of identity group – which is often a key focus in this literature. This is both in the interest of tractability and because it is not needed to explain the stylised facts we focus on.
@@ 2 Model
There are $J$ agents, with typical agent $j$. Each agent is endowed with income $w_{j}>0$ and a fixed identity $\theta_{j}$, from a finite set of possible identities $\Theta$. Agents are embedded in a weighted and directed network represented by a $J\times J$ non-negative matrix $G$. We say that agent $k$ is a neighbour of $j$ if $G_{jk}>0$, and assume that $j$ is not her own neighbour (so $G_{jj}=0$ for all $j$). Let $I(\theta)=\{k:\theta_{k}=\theta\}$ be the set of agents who have identity $\theta$.
@@
Agents simultaneously choose consumption of a visible good, $x_{j}\geq 0$.
@@
Each agent $j$ has a reference point, $R_{j}=\sum_{k\in I(\theta_{j})}G_{jk}x_{k}$; a weighted sum of the consumption of her neighbours who share her identity.
@@
Agent $j$ gets status benefits that are increasing in her own consumption, $x_{j}$, and in the average consumption of agents with the same identity as her, $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}$. And we assume that the status from each of these sources are imperfect substitutes: the marginal benefit of each is decreasing in the level the other. [^8] We also assume that agent $j$ ignores her own impact on $Y(\theta_{j})$ when choosing her consumption. This has a vanishingly small impact as $J$ grows large.
Agent $j$ also experiences dissonance costs when her consumption differs from her reference point, and these costs are convex. Note that $j$’s status benefits depend on everyone who shares her identity, but her reference point depends only on her neighbours who share her identity. There is also a convex cost of own consumption, with both absolute and marginal cost decreasing in own income. This cost function captures spending on other goods (outside of our model) in a reduced form way; implicitly assuming concave benefits from consuming the unmodelled goods.
In order to fix ideas and to show the results more clearly, we impose a particular functional form. We assume agents have the following preferences:
@@
[…]
@@
Each agent knows her own preferences, the average consumption who share her identity, and the consumption choices of her neighbours. All agents move simultaneously. So we look for Nash Equilibrium of the game.

### v7

@@ Abstract
This paper adapts ideas from social identity theory to set out a new framework for modelling conspicuous consumption. Agents derive utility from their consumption of a status good and from belonging to an identity group with high status good consumption. Importantly, these two sources of utility are substitutes. Agents also feel pressure to conform with their neighbours in a network. This framework can rationalise a set of seemingly conflicting stylised facts about conspicuous consumption that are currently explained by different families of models. In addition, our model delivers new testable predictions regarding the effect of network structure and income inequality on conspicuous consumption.
@@
Keywords:    Social identity, social networks, interpersonal comparisons, norms,
discrimination, centrality, income inequality, Keeping Up with the Joneses
Economists have long acknowledged that some goods are consumed as much for the status they convey as for their actual usefulness, a phenomenon Veblen termed conspicuous consumption. For these status goods, comparisons with others is understood to be a particularly important consideration in people’s consumption decisions. To date, an impressive body of empirical work has documented a range of stylised facts regarding how people’s consumption of status goods responds to the choices and incomes of others.
First, an influential paper by Charles et al. (2009) documents that, holding their own permanent income fixed, people in the U.S. consume less status goods (which they call ‘visible goods’) when average income of those of the same race and who live in the same state is higher. Khamis et al. (2012) find similar behaviour for caste/religious groups in India (Fact 1). But second, people consume more when average income of their geographic neighbours increases (Kuhn et al. (2011), Fact 2). And they consume more when the income of other social groups rises (Bellet and Colson-Sihra (2018), Fact 3). Additionally, people’s status good consumption is more sensitive to the incomes of others when they have higher income themselves (Kaus (2013), Fact 4). [^2] A large part of the contribution in Kaus (2013) is to extend the analysis in Charles et al. (2009) to a lower income country – South Africa. He replicates the findings for Black South Africans (a large majority of the South African population), but not for White South Africans. But the finding regarding sensitivity is novel to Kaus (2013).
Beyond these effects of changes in the level of income, Charles et al. (2009) documents that the impact of inequality can differ markedly across different groups (Fact 5). They find that higher dispersion of income within a person’s reference group leads to lower status good consumption among Whites, but higher status good consumption among Blacks. Finally, there is widespread evidence that, conditional on income, some social groups spend more on status goods than others (Charles et al. (2009); Kaus (2013); Bellet and Colson-Sihra (2025), Fact 6).
These six facts are by no means an exhaustive description of how people consume status goods. Nevertheless, we contend that they capture many of the most important features of status good consumption that have been documented empirically. Our paper presents a simple model of social identity that explains these six stylised facts as a single phenomenon. This stands in contrast to existing theory, which struggles to rationalize these behaviours as a coherent whole.
In our model, people belong to an identity group and derive utility from two sources: (1) directly from consuming status goods themselves, [^3] Own consumption can bring both intrinsic utility and individual status. and (2) from belonging to an identity with high average status good consumption, relative to that of other groups. Importantly, we assume there is some imperfect substitutability between these two sources.
Belonging to an identity group also makes people feel pressure to conform to some reference point of status good consumption for their identity group (a ‘benchmark for normal’). We assume this reference point depends only on a person’s neighbours in a social network, rather than their entire identity group. These two features – deriving benefits from group status and the pressure to conform – are both core components of social identity theory (Costa-Font and Cowell, 2015).
Our model sees agents gain utility directly from consuming the status good. This interpretation of status is the view of many sociologists (Coleman, 1994) and of some economists O’Cass and McEwen (2004) and (Becker et al., 2005). [^4] The origin of the desire for status, including evolutionary pressures, are explored in Anderson et al. (2015). Within economics, it has proved especially popular in network economics (e.g. Immorlica et al. (2017); Langtry (2023); Ghiglino and Goyal (2010)). But overall, a mainstream view in economics is that status comes only from income or wealth – and that status goods are simply a costly signal (Bagwell and Bernheim, 1996; Glazer and Konrad, 1996; Charles et al., 2009).
Note that most of the literature uses status consumption and conspicuous consumption interchangeably (see discussion in O’Cass and McEwen (2004)). We adopt the view that conspicuous consumption, that is, consumption of a good in excess of its usefulness, is the result of the good providing status and prestige.
Agents also gain status from belonging to a group with high-status members. Here, this is one with high status good consumption. But group status is relative – it depends on the group’s average consumption relative to that of the other group(s). This idea that status is relative can be traced back at least to Max Weber’s (1918) original work on status groups. It remains a firm feature of thinking about status – regardless of whether one adopts the direct status view or the signalling view. [^5] In Appendix D, we consider the alternative case where status is absolute. That is, where it depends only on the consumption of an agent’s own group.
@@
On the way to explaining the six stylised facts set out above, we provide a simple characterisation of behaviour. Equilibrium consumption of status goods depends critically on a generalised version of Bonacich centrality (Katz, 1953; Bonacich, 1987) – a well-known measure of network centrality. But here, both an agent’s own centrality and the average centrality of each identity group are important. An agent’s status good consumption increases linearly in her own centrality (a feature common to many networks models). But it also decreases in the average centrality of her own identity group and increases in the average centrality of other identity groups. In line with existing work, centrality is increasing in the strength of network connections, in an agent’s own income and in the income of those who share her identity.
The contrasting impacts of own centrality and group-average centrality is driven by the two sources of status in our model – own consumption and group-average consumption – and the fact that they are imperfect substitutes. Higher own centrality raises an agent’s status good consumption directly. But higher centrality for others in her group raises their consumption – which raises group-average consumption, in turn raising the group’s status. This higher group status reduces the marginal benefit of an agent’s own consumption, inducing her to consume less. Because group status is inherently relative, higher centrality in other groups has the opposite effect. Higher centrality, and hence higher consumption, in other groups lowers the agent’s group status. In turn, this pushes her to consume more status goods herself.
It is precisely these contrasting impacts that explain the first four stylised facts. When incomes for some agents in a given identity group rise, they consume more status goods – the traditional direct response to an increase in own income. This then creates two spillover effects. First, it increases centralities for others in that identity group. This effect pushes status good consumption of those others in the identity group up. Second, it increases average consumption of that identity group – which raises group status for that identity group. In turn, this pushes others within the identity group to consume less status goods – because it reduces the marginal benefit of an agent’s own status good consumption.
@@
Which of the two spillover effects wins out depends on how closely connected an agent $i$ is to those whose incomes increased. When those whose incomes increased are close neighbours (as in Kuhn et al. (2011) and Khamis et al. (2012)) we should expect the first effect to win out – so $i$ consumes more status goods (Fact 2). In contrast, when $i$ is not closely connected to those whose incomes increased (as in Charles et al. (2009)) we should expect the second effect to win out – delivering the seemingly counter-intuitive result that she consumes fewer status goods (Fact 1).
Notice that for other identity groups, only the second effect operates. There is no effect on own centrality in different identity groups. But higher average consumption in one identity group reduces the group status in all other groups. In turn, this pushes others in different identity groups to consume more status goods (Fact 3). The explanation for Fact 4 is even more direct. Centrality acts multiplicatively on all other terms in our characterisation of equilibrium status good consumption. And centrality is increasing in own income. So higher own incomes makes status good consumption more sensitive to changes in other primitives – including others’ incomes.
We then study the effects of changing inequality. Here, we consider a setting with a number of separate communities and homogenous incomes within each community. The impact of redistributing income across different communities depends critically on the relative densities of network connections within the affected communities. This highlights the role that network connections play in mediating the impact of inequality. Here, we present some stylised evidence of a non-monotonic relationship between income and network densities. This relationship, coupled with well-documented differences in average incomes across different groups, delivers sharply differing effects of inequality on different identity groups (Fact 5).
Finally, we extend the model to incorporate some exogenous component to group status – which we call prestige. This captures the idea that many other factors beyond status good consumption may influence an identity group’s status. One important determinant of a group’s prestige may be discrimination it faces within a society (Oldmeadow and Fiske, 2010; Fox et al., 2015). In our model, groups with lower prestige consume more status goods (all else equal). This is because agents in these groups use status good consumption to compensate for lower prestige. Differences in discrimination/prestige directly induce differences in status good consumption across groups (Fact 6). And higher consumption of status goods would lead to lower consumption of other (unmodelled) goods – including capital goods. So our model draws a novel link between discrimination and lower consumption of capital goods.
@@
The rest of the paper is organised in the usual way. Section 1 briefly reviews related literature. Section 2 presents the model. Section 3 presents the main results. Section 4 shows how our model explains the stylised facts. Then Section 5 examines the impact of inequality. Section 6 extends the model to include an exogenous component of group status, and discusses the effects of discrimination. Section 7 concludes. All proofs are in Appendix A.
@@ 1 Literature
Our paper sits within a large literature on interpersonal comparisons and conspicuous consumption. [^6] It spans disciplines, including marketing, sociology, social psychology (see Dubois and Ordabayeva (2017) and Dubois et al. (2021) for recent reviews), and economics (where it dates at least to Veblen (1899), with other early contributions from Duesenberry (1949); Festinger (1954); Frank (1985b, a) and Van de Stadt et al. (1985)). Bursztyn and Jensen (2017) provide an interesting survey of the wider topic of social pressure. Within economics, there are two broad views on why people engage in conspicuous consumption. The first assumes that people care about their income or wealth relative to others – and conspicuous consumption is used as a signal of high income/wealth. Charles et al. (2009) and Khamis et al. (2012) use this signalling approach to explain the first stylised fact (Fact 1) – when others get richer, a given agent moves down the income ranking, and so has fewer people she needs to signal she is richer than (their model follows Glazer and Konrad (1996) very closely). However, with no explicit notion of ‘neighbours’, these models struggle to explain the second stylised fact – that people consume more when their neighbours’ incomes rise. Additionally, because agents are consuming to differentiate themselves from people poorer than themselves, everyone’s consumption decisions can be sensitive to incomes in the lowest part of the distribution. Our model can speak to Fact 1 without this sensitivity.
@@
The second view assumes that people make social comparisons directly over consumption – so people gain status from conspicuous consumption itself. The core idea is that people like high consumption relative to others, not just in absolute terms – often called a “Keeping up with the Joneses” motive. Earlier work assumed people either care about their consumption relative to the population average, or to their rank within the population. Important contributions to these two approaches include Abel (1990); Clark and Oswald (1996); Ljungqvist and Uhlig (2000), and (Frank, 1985a, b, 2001; Hopkins and Kornienko, 2004) respectively. [^7] More recent contributions include Luttmer (2005); Heffetz (2011); Frank et al. (2014); Drechsel-Grau and Schmid (2014); Alvarez-Cuadrado et al. (2016); Jinkins (2016); Bertrand and Morse (2016),De Giorgi et al. (2020) and Hopkins (2024) Our paper is most closely related to a recent strand within this literature, started by Ghiglino and Goyal (2010), that uses explicit networks to allow for very rich patterns of social comparisons (more recent contributions to this strand include Immorlica et al. (2017); Ushchev and Zenou (2020); Langtry (2023) and Bramoullé and Ghiglino (2022)). These ‘Keeping up with the Joneses’ models successfully explain the second stylised fact. But they cannot explain the first. This is because they embed a desire to ‘Keep up with the Joneses’, which makes agents’ consumption choices strategic complements – an increase in some people’s income must (weakly) increase everyone’s consumption.
@@
Our paper explains a set of so-far difficult to reconcile stylised facts in a single framework. It does so by building a new type of model – one motivated by social identity, rather than signalling or ‘Keeping up with the Joneses’ motives. In doing so, we also relate to a strand of literature that examines the link between inequality and conspicuous consumption. By and large, formal models provide ambiguous results on the nature of the link. In Charles et al. (2009), a regressive redistribution of income (from a poorer agent to a richer one) increases total conspicuous consumption only if consumption is convex in income – but this is not a consistent property of observed consumption data (Heffetz, 2011). [^8] Other versions of the signalling model make similar predictions (Ireland, 2001; Heffetz, 2011; Moav and Neeman, 2012; Jinkins, 2016).
The present model assumes cardinal utilities. Bilancini and Boncinelli (2012) show that the relationship between conspicuous consumption and inequality can be sensitive to the choice of whether to use ordinal or cardinal utility. Hopkins (2024), who extends Hopkins and Kornienko (2004) to cardinal games, reaches a similar conclusion and shows that cardinal games are more general and offer a better description of the evidence.
@@
As in this literature, we do not present an ironclad relationship between inequality and conspicuous consumption. But we do provide novel predictions regarding how the nature of the relationship depends on properties of the network. Notably, it allows us to explain our fifth stylised fact – that higher dispersion in income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks. It does so without an appeal to differences in underlying preferences across groups, but instead relies on well-documented differences in average income.
@@
Finally, our paper relates to the burgeoning literature on social identity theory within economics (see Costa-Font and Cowell (2015) for a survey). The core idea in these models is that agents are part of an ‘identity’ group, and gain some status from belonging to the group, but also bear some costs from deviating from some benchmark for that group (either real differences, or in the agents’ own perception of matters). This setup was formalised in seminal contributions from Akerlof and Kranton (2000) and Shayo (2009). [^9] A number of recent papers extend or apply Shayo’s approach; see for example, Klor and Shayo (2010); Bonomi et al. (2021); Grossman and Helpman (2021); Lindqvist and Östling (2013); Atkin et al. (2021), and Shayo (2020). Ghiglino and Tabasso (2024) consider how network connections affect the choice of identity (in contrast, our paper treats identity as fixed and considers how the network affects consumption choices).
@@
Our main contribution here is to show how models of social identity can help resolve a puzzle regarding conspicuous consumption behaviour – something these models have not been used for to date. However, we abstract away from the choice of identity group – which is often a key focus in this literature. This is both in the interest of tractability and because it is not needed to explain the stylised facts we focus on. From a more technical standpoint, aside from this, our utility formulation closely follows Shayo (2009). However, departing from most of the literature, but closer to Atkin et al. (2021), we make the benchmark (often called ‘prototype’) as well as status both endogenous.
@@ 2 Model
There are $J$ agents, with typical agent $j$. Each agent is endowed with real income $w_{j}\in[1,\infty)$ and a fixed identity, from two possible identities: $\theta_{j}\in\{A,B\}$. Agents are embedded in a weighted and directed network represented by a $J\times J$ non-negative matrix $G$. We say that agent $k$ is a neighbour of $j$ if $G_{jk}>0$, and assume that $j$ is not her own neighbour (so $G_{jj}=0$ for all $j$). Let $I(\theta)=\{k:\theta_{k}=\theta\}$ be the set of agents who have identity $\theta$.
@@
Agents simultaneously choose consumption of a status good, $x_{j}\geq 0$.
@@
Each agent $j$ has a reference point, $R_{j}=\alpha_{j}\sum_{k\in I(\theta_{j})}g_{jk}x_{k}$, where $g_{jk}=\frac{G_{jk}}{\alpha_{j}}$ and $\alpha_{j}=\sum_{k}G_{jk}$; a weighted average of the status good consumption of her neighbours who share her identity.
@@
The agent $j$ receives utility benefits from her own consumption, $x_{j}$, and from the status of the identity group she identifies with [^10] The assumption that status goods have both intrinsic and status benefits is also present in Immorlica et al. (2017).. The status term is given by the average consumption of her identity group relative to the other identity group. For simplicity, we model this relative term as the ratio of average consumption of each group: $Y_{\theta}=\frac{\mathbf{x}_{\theta}}{\mathbf{x}_{-\theta}}$, where $\mathbf{x}_{\theta}=\frac{1}{|I(\theta)|}\sum_{k\in I(\theta)}x_{k}$ for $\theta\in\{A,B\}$. Importantly, we assume that own consumption and identity status are imperfect substitutes: the marginal benefit of each is decreasing in the level of the other.
Agent $j$ also experiences dissonance costs when her consumption differs from her reference point, and these costs are convex. This is standard in identity theory. Note that $j$’s status benefits depend on everyone who shares her identity, but her reference point depends only on her neighbours who share her identity.
Finally, there is also a convex cost of own consumption, with both absolute and marginal cost decreasing in own income. This cost function captures spending on other goods (outside of our model) in a reduced form way; implicitly assuming concave benefits from consuming the unmodelled goods (see discussion in Section 2.2). We assume agents have the following preferences:
@@
with $\beta>0$ and $S^{\prime}_{x}>0,S^{\prime}_{Y}>0,S^{\prime\prime}_{xY}<0$, for $x\in[0,X]$ and $Y\in\left[\frac{1}{Z},Z\right]$ for some $X>0,Z\in(0,\infty)$. We provide a sufficient characterisation of $X$ and $Z$ in Section 2.1. In the interest of tractability, we assume that the utility from own consumption is linear in $x_{j}$, so $S^{\prime}_{x}$ is not a function of $x_{j}$. A simple example of a function that meets these assumptions is $S(x,Y)=\eta\,x+Y-\gamma\,x\,Y$, where $\eta>0$ and $\gamma<\min\{\frac{\alpha}{Z},\frac{1}{X}\}$.
For tractability we also assume that agent $j$ ignores her own impact on $Y_{j}$ when choosing her consumption. When the number of agents is large, $j$’s impact on $Y_{j}$ is small, so this assumption has less and less ‘bite’ as $J$ grows. An alternative approach would be to assume an infinite number of agents, so $j$’s impact on $Y_{j}$ is exactly zero. Such an approach would require extra notation without adding insight.
@@
All of the primitives – the network, incomes, functional forms, the value of beta – are common knowledge. All agents move simultaneously. So we look for Nash Equilibrium of the game.